# Classical State Detection Using Quantum State Tomography

### v1

@@ Abstract
In this paper, we inject a weak coherent light with a well-defined polarization state into a idler photon channel, which is generated from a fiber-based entangled photon source. We then perform a full quantum state tomography on the classically mixed idler photon and its paired signal photon. The weak coherent light is interpreted as a measurement apparatus, which introduces noise photons and induces a classical state at the same optical wavelength as the idler photon. This classical state corresponds to the polarization state of the injected weak coherent light. To extract information about the classical state, we assume the initial state is a Werner stateWerner_pra_1989 ; Munro_pra_2001 and the final state is a $X$- stateDonnert_nmeth_2007 ; Luo_pra_2008 . We model the final state as a combination of its quantum part and a classical-quantum (CQ) state. The CQ state arises from a local quantum operation (measurement) on the idler photon. We numerically extract the density matrix of the classical state from the CQ state by solving the diagonal and off-diagonal elements of the measured final state’s density matrix with the theoretical model. The resulting classical state’s density matrix corresponds to the local quantum operation acting on the idler photon. The detected classical state agrees with the experimental result. We then evaluate the diagonal quantum discord of the states involved in this study. Our experimental approach demonstrates both the practicality and robustness of detecting classical state by using quantum state tomography. This method could significantly influence the development of tools for classical-quantum coexistence in networking techniques. Furthermore, our algorithm may prove highly beneficial for advancing quantum key distribution protocols that involve the coexistence of the weak coherent statesBarbosa_2003 ; Yuen_2003 ; Decoy_2005 ; Sua_2011 and entangled photon statesBB_2014 ; BB_1992 ; Gisin_2002 .
@@
Figure 1: A counter-propagating scheme is used to generate two-photon polarization entangled photon pairs. The polarization state of the weak coherent light is prepared in free-space and then mixed with the idler photon through a DWDM. The classically mixed idler photon and the signal photon are analyzed by two spatially separate polarization analyzers ($\rm{PA_{s,i}}$). HWP; Half-wave plate. QWP; Quarter-wave plate. PBS; polarization beam splitter. SPD; Single photon detector.
We first prepare the classical state $H$ and mix with the idler photon. We use two single-photon detectors (NuCrypt CPDS-4), operating at 50 MHz, to measure coincidences and accidental-coincidences between the idler and signal photons. A total of 16 settings are used for the HWPs and QWPs in the $\rm{PA_{s}}$ and $\rm{PA_{i}}$. For each setting, we measure the coincidence counts (CC) and accidental-coincidence (AC) with an integration time of 2 seconds, corresponding to 100M sampling gates. Using the maximum likelihood estimation, we reconstruct the $4\times 4$ density matrix of the final state $\rho^{f}_{X}$ in the HV basis. We repeat the state tomography for different classical states, $A$ and $L$, mixed with the idler photon. We assume the final state is an $X$-state with a small admixture of a classical-quantum state. Let the experimentally measured final state be denoted as $\rho^{f}_{X}$. We then define the numerically modeled final state $\rho^{n}_{X}$ as
@@
where $\rho_{X}(C)$ represents the quantum part of the measured $\rho^{f}_{X}$ state, and the $C=\{H,A,L\}$ denotes the injected classical state. The state $\rho_{w}(C)$ is a Werner state, which is shared by the photon pairs before mixing the classical light in the idler channel. We obtain the Werner state $\rho_{w}(C)$ by subtracting the accidental-coincidences from the coincidences in the state tomography measurement. We could measure the Werner state by blocking the weak coherent light. However, we avoid performing two separate QST measurements on this system. The term $2(\mathcal{E}\rho_{w}(C)\mathcal{E}^{\dagger})$ is the classical-quantum state. The parameter $x$ denotes the small contribution from the classical-quantum state. The operator $\mathcal{E}=\mathcal{I}\bigotimes\mathcal{B}$ describes the local measurements $\mathcal{B}$ on the idler photon, where $\mathcal{I}$ is the identity operator acting on the signal photon. In our scenario, the weak coherent light increases the accidental-coincidences and induces the classical state, which can be interpreted as local measurement $\mathcal{B}$ on the idler photon. Therefore, the operator $\mathcal{B}$ is a $2\times 2$ density matrix representing the classical state $C$. The quantum discord for the classical-quantum state is zero. We determine $\rho_{X}(C)$ and the parameter $x$ by solving the diagonal and off-diagonal density elements of the measured density matrix $\rho^{f}_{X}$ with those of $\rho^{n}_{X}(C)$, as defined in Eq.(1). There exists a unique solution in solving Eq.(1) that determines the operator $\mathcal{B}$. The resulting $\mathcal{B}$ is consistent with the classical state $C$ prepared by the weak coherent light.
@@
where $c_{1,2,3}$ are the real parameters. The Werner state is obtained by setting $c_{3}=p$, $c_{1}=p$, and $c_{2}=-p$, where $p$ is the probability of the maximally entangled state. In this work, we consider the $X$-state $\rho_{X}$ of the form
@@
The density matrix elements of $\rho^{n}_{X}(C)$ in Eq.(1) are complex functions of the parameters $p$, $x$, and $a$. For simplicity, we use experimental data to demonstrate the algorithm used to solve Eq.(1).
II.1 2.1 Detection of Classical State H
@@
We generate a two-photon polarization-entangled state from the CPS and perform the QST with the weak coherent light in the idler channel. We reconstruct the density matrix of the final state $\rho^{f}_{X}(H)$ (see the Methods for the measured density matrix of the classically mixed idler and signal photons as shown in Eq.(15)). For solving Eq.(1), we use the measured density matrix elements of the final state, $\rho^{f}_{X}(H)$ : $p^{f}_{11}$, $p^{f}_{14}$, $p^{f}_{23}$ and $p^{f}_{13}$.
@@
The $X$-state density matrix elements of $\rho^{n}_{X}(H)$ are comparable to those of the measured final state $\rho^{f}_{X}(H)$ as shown in Eq.(15) in the Methods section. We obtain the parameters $x$ =0.0334, cc=0.5997, and aa=-0.0081. The probability of $\rho_{X}(H)$ in Eq.(1) is given by $c\times p_{w}=0.5997\times 0.972=0.582$. This relatively low probability results from mixing the weak coherent light with the idler photon, which increases accidental-coincidences in the state tomography measurement. The parameter $x$ is only $3.0\%$, indicating a small contribution from the classical-quantum state to the final state $\rho^{n}_{X}(H)$. However, this contribution is sufficient to detect the classical state $H$ in the idler channel. If we choose other operators $\mathcal{B}=(V,D,A,R,L)$ to compute $2(\mathcal{E}\rho_{w}(H)\mathcal{E}^{\dagger})$ in Eq.(1), the Eq.(1) cannot be solved, and the parameter $x$ cannot be obtained. Hence, the classical state $H$ is the only solution available for solving the final state $\rho^{n}_{X}(H)$. Since the parameter $x$ is small, the $\rho^{n}_{X}(H)$ is primarily contributed from its quantum part, $\rho_{X}(H)$.
II.2 2.2 Detection of Classical State A
Using a similar approach to that in the previous section, we inject the classical state $A$ into the idler channel. We observe that the single counts of the idler channel increase by a factor of 10, and the CAR decreases from 4 to 1. We reconstruct the density matrix of the measured final state (see the Methods for the measured density matrix $\rho^{f}_{X}(A)$ in Eq.(16)). The density matrix elements $p^{f}_{11}$, $p^{f}_{14}$, and $p^{f}_{23}$, along with the condition $p^{f}_{12}<0$ from the measured final state $\rho^{f}_{X}(A)$, are used to solve Eq.(1). The accidental-subtracted Werner state $\rho_{w}(A)$ is obtained as;
@@
where the imaginary part of $\rho_{w}(A)$ can be neglected. The probability of the Werner state, $p_{w}$, is 0.937. This is lower than in the previous case because more classical photons are injected in the idler channel. The state $2(\mathcal{E}\rho_{w}(A)\mathcal{E}^{\dagger})$ is then computed as:
@@
II.3 2.3 Detection of Classical State L
@@
We find the probability of the Werner state to be $p_{w}=0.909$, which is expected due to the increased number of classical photons injected in the idler channel and the extinction ratio of the PBS in the $PA_{i}$. We then compute $2(\mathcal{E}\rho_{w}(L)\mathcal{E}^{\dagger})$ as follow:
@@
II.4 2.4 The relative strength between $p^{f}_{11}$ and $p^{f}_{14}$
@@
The existence of a viable solution to Eq.(1) in our algorithm relies on the properties of the $X$- state and the off-diagonal elements of the classical-quantum state. However, the relative strength between $p^{f}_{11}$ and $p^{f}_{14}$, along with the sign of the parameter $a$ (i.e., whether $a>0\,\rm{or}\,a<0$) in the quantum part of the measured final state, are critical constraints for obtaining such a solution. A solution to Eq.(1) may not be attainable if the measured final state does not satisfy these constraints, as discussed in previous sections.
Finally, our algorithm can be incorporated into quantum process tomographyChuang_JMO_1997 ; Altepeter_PRL_2003 to detect quantum processes induced by the environment. Quantum processes or operators can be extracted from the process tomography, including complete projection measurements (i.e., von Neumann measurements) expressed in terms of the Pauli matrices ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$) and the identity.

### v2

@@ Abstract
In this paper, we inject a weak coherent light with a well-defined polarization state into a idler photon channel, which is generated from a fiber-based entangled photon source. We then perform a full quantum state tomography on the classically mixed idler photon and its paired signal photon. The weak coherent light is interpreted as a measurement apparatus, which introduces noise photons and induces a classical state at the same optical wavelength as the idler photon. This classical state corresponds to the polarization state of the injected weak coherent light. To extract information about the classical state, we assume the initial state is a Werner stateWerner_pra_1989 ; Munro_pra_2001 and the final state is a $X$- stateDonnert_nmeth_2007 ; Luo_pra_2008 . We model the final state as a combination of its quantum component and a classical-quantum (CQ) state. The CQ state arises from a local quantum operation (measurement) on the idler photon. We numerically extract the density matrix of the classical state from the CQ state by solving the diagonal and off-diagonal elements of the measured final state’s density matrix with the theoretical model. The resulting classical state’s density matrix corresponds to the local quantum operation acting on the idler photon. The detected classical state agrees with the experimental result. We then evaluate the diagonal quantum discord of the states involved in this study. Our experimental approach demonstrates both the practicality and robustness of detecting classical state by using quantum state tomography. This method could significantly influence the development of tools for classical-quantum coexistence in networking techniques. Furthermore, our algorithm may prove highly beneficial for advancing quantum key distribution protocols that involve the coexistence of the weak coherent statesBarbosa_2003 ; Yuen_2003 ; Decoy_2005 ; Sua_2011 and entangled photon statesBB_2014 ; BB_1992 ; Gisin_2002 .
@@
Figure 1: A counter-propagating scheme is used to generate two-photon polarization entangled photon pairs. The polarization state of the weak coherent light is prepared in free-space and then mixed with the idler photon through a DWDM. Transmission from port R to port C of the DWDM provides 23 dB of attenuation for weak coherent light at 1561 nm. The classically mixed idler photon and the signal photon are analyzed by two spatially separate polarization analyzers ($\rm{PA_{s,i}}$). HWP; Half-wave plate. QWP; Quarter-wave plate. PBS; polarization beam splitter. SPD; Single photon detector.
We first prepare the classical state $H$ and mix with the idler photon. We use two single-photon detectors (NuCrypt CPDS-4), operating at 50 MHz, to measure coincidences and accidental-coincidences between the idler and signal photons. A total of 16 settings are used for the HWPs and QWPs in the $\rm{PA_{s}}$ and $\rm{PA_{i}}$. For each setting, we measure the coincidence counts (CC) and accidental-coincidence (AC) with an integration time of 2 seconds, corresponding to 100M sampling gates. Using the maximum likelihood estimation, we reconstruct the $4\times 4$ density matrix of the final state $\rho^{f}_{X}$ in the HV basis. We repeat the state tomography for different classical states, $A$ and $L$, mixed with the idler photon. We then define the numerically modeled final state $\rho^{n}_{X}$ as
@@
where $\rho_{X}(C)$ represents the quantum component of the measured $\rho^{f}_{X}$ state, and the $C=\{H,A,L\}$ denotes the injected classical state. The state $\rho_{w}(C)$ is a Werner state, which is shared by the photon pairs before mixing the classical light in the idler channel. We obtain the Werner state $\rho_{w}(C)$ by subtracting the accidental-coincidences from the coincidences in the state tomography measurement. In our detection system, a coincidence count (CC) is recorded when both avalanche photodiodes (APDs) detect photons within the same gated time interval, whereas an accidental-coincidence count (AC) is recorded when both APDs detect photons in adjacent gated time intervals. The coincidence counts arise from both photon pairs and uncorrelated photons, while the accidental coincidence counts originate solely from uncorrelated photons. These uncorrelated photons mainly come from: (i) Raman photons generated by spontaneous Raman scattering in the optical fiber; (ii) loss of one photon from a pair in either the signal or idler channel; and (iii) injected classical light in the idler channel. By subtracting the accidental-coincidences from the coincidence counts (i.e., $\rm{CC-AC}$), we obtain the true coincidence counts arising from photon pairs. We use these true coincidence counts from the 16 measurement settings of the half-wave plate (HWP) and quarter-wave plate (QWP) in the quantum state tomography (QST) to reconstruct the Werner state. We could measure the Werner state by blocking the weak coherent light. However, we avoid performing two separate QST measurements on this system. The reason is the accidental-subtracted Werner state has a probability identical to that of the Werner state measured when the weak coherent light is blocked. The term $2(\mathcal{E}\rho_{w}(C)\mathcal{E}^{\dagger})$ is the classical-quantum state. The parameter $x$ denotes the small contribution from the classical-quantum state. The operator $\mathcal{E}=I\bigotimes\mathcal{B}$ indicates that $\mathcal{B}$ acts on the idler channel, while the identity operator $I$ acts on the signal channel, since the signal photon is not mixed with the classical light. The symbol $\bigotimes$ denotes the tensor product. Consequently, the accidental-subtracted Werner state is transformed as $2(\mathcal{E}\rho_{w}(C)\mathcal{E}^{\dagger})$, resulting in a classical-quantum state, which is a $4\times 4$ matrix.
We adopt the quantum discord framework to calculate the classical correlation of an open bipartite system, in which the environment performs a projective (von Neumann) measurement on the idler channel. The operator $\mathcal{B}$ represents the outcome of the von Neumann projection on the idler channel and can be expressed as a linear combination of the Identity matrix and Pauli matrices, which are $2\times 2$ matricesLuo_pra_2008 ; Luo_2008 . When $\mathcal{B}=H$, the polarization of the idler photon is projected onto the horizontal ($H$) state. Similarly, when $\mathcal{B}=A$ and $\mathcal{B}=L$, the polarization of the idler photon is projected onto the anti-diagonal ($A$) and left-circular ($L$) polarization states, respectively. In our scenario, the weak coherent light increases the accidental-coincidences and induces the classical state, which can be interpreted as local measurement $\mathcal{B}$ on the idler photon. Therefore, the operator $\mathcal{B}$ is a $2\times 2$ density matrix representing the classical state $C$. The quantum discord for the classical-quantum state is zero. We determine $\rho_{X}(C)$ and the parameter $x$ by solving the diagonal and off-diagonal density elements of the measured density matrix $\rho^{f}_{X}$ with those of $\rho^{n}_{X}(C)$, as defined in Eq.(1). Eq.(1) admits a single admissible solution for the operator $\mathcal{B}$. Specifically, our method identifies a single polarization state from the discrete set of polarization states ($H$, $V$, $D$, AA, $R$, $L$), and does not extend to superpositions of these states. The resulting $\mathcal{B}$ is consistent with the classical state $C$ prepared by the weak coherent light.
@@
where $c_{1,2,3}$ are the real parameters. The Werner state is obtained by setting $c_{3}=p$, $c_{1}=p$, and $c_{2}=-p$. The Werner state is a mixture of the maximally entangled state $|\psi_{\circ}\rangle$ with the probability $p_{w}$ and the maximally mixed state as given by Werner_pra_1989 ; Munro_pra_2001 ,
$\rho_{w}=p_{w}|\psi_{\circ}\rangle\langle\psi_{\circ}|+\frac{1-p_{w}}{4}\mathcal{I}.$ (3)
where the $\mathcal{I}$ is the $4\times 4$ identity matrix. The Werner state can be represented as a $4\times 4$ matrix:
${\rho}_{w}=\begin{pmatrix}\frac{1+p_{w}}{4}&0&0&\frac{p_{w}}{2}\\ 0&\frac{1-p_{w}}{4}&0&0\\ 0&0&\frac{1-p_{w}}{4}&0\\ \frac{p_{w}}{2}&0&0&\frac{1+p_{w}}{4}\end{pmatrix}$ (4)
We calculate the probability $p_{w}=\sum^{n=6}_{i}p^{\prime}_{i}/n$, where $p^{\prime}_{1(2)}=4{\rho_{w}}_{11(44)}-1$, $p^{\prime}_{3(4)}=2{\rho_{w}}_{14(41)}$, and $p^{\prime}_{5(6)}=1-4{\rho_{w}}_{22(33)}$. The density matrix elements ${\rho_{w}}_{11}$, ${\rho_{w}}_{44}$, ${\rho_{w}}_{14}$, ${\rho_{w}}_{41}$, ${\rho_{w}}_{22}$, and ${\rho_{w}}_{33}$ are obtained from the accidental-subtracted Werner state. In this work, we consider the $X$-state $\rho_{X}$ of the form
@@
In general, the local measurement operator $\mathcal{B}$ is extremely weak and therefore experimentally undetectable. To simulate its action, we inject classical light with a well-defined polarization state ($H$, AA, or $L$) into the idler channel. In the following sections, we first demonstrate the validity of this approach by: (i) increasing the intensity of the classical light to approximately 1.5 times that of the idler photon, (ii) preserving a coincidence-to-accidental ratio (CAR) greater than 1 (or equivalently, $\rm{CC}-\rm{AC}>0$), and (iii) preparing a simple horizontal ($H$) polarization state. We then increase the intensity of the classical field for the $A$ and $L$ polarization states. We want to demonstrate that, as the intensity of the classical field increases, our approach remains robust against (i) experimental imperfections (i.e., errors in $\rm{CC}-\rm{AC}$ subtraction) and (ii) limitations of the experimental apparatus (i.e., the extinction ratio of the PBS in the $PA_{i}$ ). Even under these imperfect conditions, our approach is still capable of correctly identifying the corresponding classical polarization states.
Increasing the intensity of the classical light reduces the probability of the final state and its quantum component. However, as discussed above, the intensity does not affect the classical-quantum (CQ) state because the Werner state is obtained from the true coincidences. The polarization state of the classical light determines the operator $\mathcal{B}$. The parameter $p_{w}$ represents the probability of the accidental-subtracted Werner state. The probability of the quantum component $\rho_{X}$ of the final state is reduced by the factor $c$, such that it becomes $p=c\times p_{w}$. The parameter $x$, with $1>x>0$, denotes the contribution of the CQ state to the final state. The parameter $a$ is introduced in the $X$-state in Eq.(5), making it slightly different from the Werner state (which corresponds to $a=0$). The $X$-state in Eq.(5) is therefore a ”Werner-like” $X$-state introduced in this work. The density matrix elements of $\rho^{n}_{X}(C)$ in Eq.(1) are complex functions of the parameters $c$, $p_{w}$, $x$, and $a$. For simplicity, we use experimental data to demonstrate the algorithm used to solve Eq.(1).
II.1 2.1 Procedure to identify Operator $\mathcal{B}$
As mentioned above, the operator $\mathcal{B}$ used in $2(\mathcal{E}\rho_{w}(C)\mathcal{E}^{\dagger})$ in Eq.(1) corresponds to the polarization state of the injected classical light. For each choice of $\mathcal{B}$, we construct the corresponding $4\times 4$ matrix of the classical-quantum (CQ) state. For each CQ state, specific density-matrix elements (real and imaginary parts) can be identified and used as constraints (i.e., greater or less than zero) in solving Eq.(1).
The procedure used to identify the operator $\mathcal{B}$ is outlined as follows:
1. We perform quantum state tomography (QST) on the final state and reconstruct its density matrix.
2. Using the true coincidence counts from the QST data, we reconstruct the $4\times 4$ matrix of the accidental-subtracted Werner state.
3. The probability $p_{w}$ is then obtained from the accidental-subtracted Werner state.
4. We solve Eq.(1) by (i) using the density-matrix elements of the measured final state and (ii) imposing the constraints derived from each candidate $\mathcal{B}$ operator in the corresponding CQ state. Since we have no prior knowledge of the polarization state of the classical field (i.e., the operator $\mathcal{B}$), we perform Mathematica function NSolve separately for each possible $\mathcal{B}$, using the constraints obtained by its corresponding CQ state.
II.2 2.2 Detection of Classical State H
@@
We generate a two-photon polarization-entangled state from the CPS and perform the QST with the weak coherent light in the idler channel. We reconstruct the density matrix of the final state $\rho^{f}_{X}(H)$ (see the Methods for the measured density matrix of the classically mixed idler and signal photons as shown in Eq.(17)). For solving Eq.(1), we use the measured density matrix elements of the final state, $\rho^{f}_{X}(H)$ : $p^{f}_{11}$, $p^{f}_{14}$, $p^{f}_{23}$ and $p^{f}_{13}$. The measured final state has the same structure as the $X$-state given in Eq.(5), which can ideally be described by two parameters, $p$ and $a$. However, the experimentally measured final state is not perfect; for example, the density matrix elements $p^{f}_{11}$ and $p^{f}_{44}$ are not equal. We therefore select $p^{f}_{11}$, $p^{f}_{14}$, $p^{f}_{23}$ and $p^{f}_{13}$ from the measured final state to solve the Eq.(1). This approach allows us to obtain a numerically solved final state whose density matrix elements closely match those of the measured final state, i.e., $p^{n}_{44}\sim p^{f}_{44}$, etc. If $p^{f}_{44}$ is included in addition to the four selected density matrix elements, Eq.(1) cannot be solved because the experimentally measured values $p^{f}_{11}$ and $p^{f}_{44}$ are not equal. Therefore, choosing these four density matrix elements provides the optimal and consistent set of constraints for solving Eq.(1).
@@
The $X$-state density matrix elements of $\rho^{n}_{X}(H)$ are comparable to those of the measured final state $\rho^{f}_{X}(H)$ as shown in Eq.(17) in the Methods section. We obtain the parameters $x$ =0.0334, cc=0.5997, and aa=-0.0081. The probability of $\rho_{X}(H)$ in Eq.(1) is given by $c\times p_{w}=0.5997\times 0.972=0.582$. This relatively low probability results from mixing the weak coherent light with the idler photon, which increases accidental-coincidences in the state tomography measurement. The parameter $x$ is only $3.0\%$, indicating a small contribution from the classical-quantum state to the final state $\rho^{n}_{X}(H)$. However, this contribution is sufficient to detect the classical state $H$ in the idler channel. If we choose other operators $\mathcal{B}=(V,D,A,R,L)$ to compute $2(\mathcal{E}\rho_{w}(H)\mathcal{E}^{\dagger})$ in Eq.(1), the Eq.(1) cannot be solved, and the parameter $x$ cannot be obtained. For example, if we assume $\mathcal{B}=A$ and impose the condition $p^{n}_{\rm{off\,diag.}}$ $p^{f}_{\rm{off\,diag.}}<0$ (i.e., $p^{f}_{12}$, $p^{f}_{21}$, $p^{f}_{13}$, $p^{f}_{31}$, $p^{f}_{24}$, $p^{f}_{42}$, $p^{f}_{34}$, $p^{f}_{43}$)<0)<0 (as discussed in Section 2.3), Eq.(1) has no solution. This is because the experimentally measured off-diagonal elements $p^{f}_{\rm{off\,diag.}}$ of the final state (see Eq.(17) in Section 5.1) are not all negative. Similarly, for $\mathcal{B}=V$, $L$, and other polarization states, Eq.(1) has no solution, whereas a consistent solution exists only for $\mathcal{B}=H$. This procedure therefore allows us to identify the polarization state of the injected classical light. Since the parameter $x$ is small, the $\rho^{n}_{X}(H)$ is primarily contributed from its quantum part, $\rho_{X}(H)$.
II.3 2.3 Detection of Classical State A
Using a similar approach to that in the previous section, we inject the classical state $A$ into the idler channel. We observe that the single counts of the idler channel increase by a factor of 10, and the CAR decreases from 4 to 1. We reconstruct the density matrix of the measured final state (see the Methods for the measured density matrix $\rho^{f}_{X}(A)$ in Eq.(18)). The accidental-subtracted Werner state $\rho_{w}(A)$ is obtained as;
@@
where the imaginary part of $\rho_{w}(A)$ can be neglected. The probability of the Werner state, $p_{w}$, is 0.937. This is lower than in the previous case because more classical photons are injected in the idler channel and imperfect $\rm{CC-AC}$ subtraction. The state $2(\mathcal{E}\rho_{w}(A)\mathcal{E}^{\dagger})$ is then computed as:
@@
II.4 2.4 Detection of Classical State L
@@
We find the probability of the Werner state to be $p_{w}=0.909$, which is expected due to the increased number of classical photons injected in the idler channel, imperfection in subtracting $\rm{CC-AC}$, and the extinction ratio of the PBS in the $PA_{i}$. We then compute $2(\mathcal{E}\rho_{w}(L)\mathcal{E}^{\dagger})$ as follow:
@@
[…]
II.5 2.5 The relative strength between $p^{f}_{11}$ and $p^{f}_{14}$
@@
The existence of a viable solution to Eq.(1) in our algorithm relies on the properties of the $X$- state and the off-diagonal elements of the classical-quantum state. However, the relative strength between $p^{f}_{11}$ and $p^{f}_{14}$, along with the sign of the parameter $a$ (i.e., whether $a>0\,\rm{or}\,a<0$) in the quantum component of the measured final state, are critical constraints for obtaining such a solution. A solution to Eq.(1) may not be attainable if the measured final state does not satisfy these constraints, as discussed in previous sections.
In the classical-quantum coexistence schemeGamze_2025 , classical light co-propagates with quantum light in an optical fiber, but at different optical wavelengths. The classical light can be used to monitor and compensate for the polarization drift experienced by the quantum lightGamze_Q2025 . The classical and quantum signals are separated by inserting a wavelength-division multiplexer (WDM) immediately before the polarization analyzer and single-photon detectors. To achieve effective polarization compensation for the quantum light, the wavelengths of the classical and quantum signals are ideally matched. In a recent polarization-entangled photon distribution experiment over New York City fiberMehdi_JOC_2025 ; Mehdi_PRX_2025 , classical light at the same wavelength as the quantum light was used to perform automated polarization compensation. However, during this compensation process, the quantum light was turned off. In contrast, our method enables measurement of the polarization state of the classical light using QST, with both the classical and quantum signals passing through the polarization analyzer and into single-photon detector. We believe that this approach can improve polarization compensation in classical-quantum coexistence networks, such as quantum wrapping, where classical and quantum signals coexist in different time domains.
Finally, our algorithm can be incorporated into machines learning Guo_2025 ; Torlai_NP_2018 and quantum process tomographyChuang_JMO_1997 ; Altepeter_PRL_2003 with unsupervised learningTorlai_NC_2023 to detect quantum processes induced by the environment. Quantum processes or operators can be extracted from the process tomography, including complete projection measurements (i.e., von Neumann measurements) expressed in terms of the Pauli matrices ($\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$) and the identity matrix.